OpenStudy (anonymous):

-11-(x+30)=-41-2(5x-18)

OpenStudy (anonymous):

-11-x-30=-41-10x+36 9x-36=0 9(x-4)=0 x=4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome